OpenStudy (angelab97):

The day of a quiz, only Michelle was absent. The average(arithmetic mean) grade of the other students was 85. When Michelle took a makeup quiz, her grade was 30, which lowered the class's average to 80. How many students are in the class? Please explain how to do this problem

OpenStudy (blurbendy):

x/c = 85 c = # students that didnt take a make up quiz x = all other scores added (x + 30) / (c + 1) = 80 x + 30 = 80(c+1) x + 30 = 80c + 80 x = 80c + 50 So, x = 85c 85c = 80c + 50 5c = 50 c = 10

OpenStudy (blurbendy):

10 + Michelle = 11 students

OpenStudy (angelab97):

thank you so much!

OpenStudy (blurbendy):

yw